x/sqrt(x^2+7)=y solve step by step for x
Answer by akmaths(3)

x/sqrt.(x^2 +7 ) = y
squaring on both sides, we have:
x^2
__________ = y^2 therefore
x^2 + 7
x^2 = x^2*y^2 + 7y^2 or
x^2 - x^2y^2 = 7y^2 or
x^2(1 -y^2) = 7y^2 or
x^2 = 7y^2 / 1 - y^2 or
x = sqrt.( 7y^2 / 1-y^2 ) ans.